There are balls packaged in twos and threes. If I have 38 packages adding up to 100 balls, how many of each package did I use?
Answer by stanbon(75887)

There are balls packaged in twos and threes. If I have 38 packages adding up to 100 balls, how many of each package did I use?
------------------
Let "x" be the number of 2-ball packages: # of balls in them = 2x
The number of 3-ball packages is "38-x": # of balls in them = 3(38-x)
-------------------------
EQUATION:
balls + balls = balls
2x + 3(38-x) = 100
-x + 114 = 100
x = 14 (# of 2-ball packages)
38-x = 24 (# of 3-ball packages)
----------------
Cheers,
Stan H.